# Reservoir-assisted quantum battery charging at finite temperatures

### v2

@@ II.2 The related performance parameter of the quantum battery
where $H_{B}=\omega_{0}\sigma_{B}^{+}\sigma_{B}^{-}$, $\rho_{B}(t)={\rm Tr}_{C}[\rho(t)]$ and $\rho_{B}(0)=|g\rangle_{B}\langle g|$ are the free Hamiltonian, the reduced density matrix and the initial state of the quantum battery, respectively. When the quantum battery is in the excited state $|e\rangle$, it means that it is fully charged, and the corresponding maximum stored energy is $E_{B}^{\rm max}=\omega_{0}$. To evaluate the useful work of the quantum battery storage, we use the concept of ergotropy Allahverdyan et al. (2004); Francica et al. (2017) and its form is as follows
@@
it describes the maximum energy that can be extracted from the battery. The second term in Eq. (9) indicates the minimum energy that cannot be extracted from the battery by executing all unitaries $U$ in the system, and the state corresponding to the minimum energy is known as the passive state Pusz and Woronowicz (1978); Lenard (1978).
@@
III THE CHARGING PROCESS OF A QUANTUM BATTERY IN DIFFERENT THERMAL RESERVOIR ENVIRONMENTS
III.1 The charging process of quantum battery in a bosonic thermal reservoir
In this part, we first consider the case where the quantum battery is in a bosonic thermal reservoir. After setting $\mathop{\dot{\rho}}=0$ in Eq. (7), the steady-state solutions of the system are obtained (see Appendix A). Combined with Eq. (8), the stored energy of the battery is finally obtained as follows
@@ III.1 The charging process of quantum battery in a bosonic thermal reservoir
This clearly shows that when $\delta=f/\Gamma_{C}=1$, the stored energy of the battery reaches its maximum value, that is, $E_{B}^{\rm max}(\infty)=\omega_{0}$. This implies that the effect of spontaneous emission on the charger can be offset by adjusting the feedback intensity $f$, enabling full charging of the quantum battery, this result is consistent with Ref. Mitchison et al. (2021). However, in practice, a quantum battery cannot be perfectly isolated from its surroundings, so it is necessary to consider placing it in an open system. To discuss the influence of feedback strength $f$ and charger dissipation on quantum battery energy storage, we conduct corresponding numerical simulations according to Eq. (12) subsequently. In the absence of special instructions, the remaining parameters are set as follows: $\eta=1$, $g=0.01\omega_{0}$, and $\Gamma_{B}=0.1g$.
@@
Figure 2: Under a bosonic thermal reservoir, the stored energy of the quantum battery is a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger. The temperature of the bosonic thermal reservoir is $T=0$ $(n_{b}=0)$ in (a), while in (b) the temperature of the bosonic thermal reservoir is $T=10$ ($n_{b}\approx 9.51$). Solid red circles represent the optimal stored energy of the quantum battery, with the corresponding optimal charging parameters being $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. Other parameters are $g=0.01\omega_{0}$ and $\Gamma_{B}=0.1g$, respectively.
@@
Figures 2 (a) and (b) show the stored energy of the quantum battery as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger when the battery is in a dissipative bosonic thermal reservoir with temperature $T=0$ ($n_{b}=0$) and $T=10$ ($n_{b}\approx 9.51$), respectively. The results show that the optimal stored energy of the battery is regulated by the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger, there exists a set of optimal parameters: $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. Here, $\delta=1$ means that the feedback strength $f$ is equal to the dissipation rate $\Gamma_{C}$ of the charger (i.e.,$f=\Gamma_{C}$).
@@
To determine whether optimal charging conditions are affected by the dissipation rate of the battery $\Gamma_{B}$, we explored the influence of the feedback parameter $\delta$ and the dissipation rate of the charger and the quantum battery on the stored energy of the quantum battery in steady state in Fig. 3. The thermal reservoir temperature for Figs. 3 (a) and (b) is $T=0$ ($n_{b}=0$) and $T=10$ ($n_{b}\approx 9.51$), respectively. The results show that when the battery dissipation rate is zero ($\Gamma_{B}=0$), there is no restriction on the magnitude of the dissipation rate $\Gamma_{C}$ of the charger, as long as the optimal feedback parameter satisfies $\delta=f/\Gamma_{C}=1$, the quantum battery can be fully charged. This is consistent with the results of the previous analysis. However, when the quantum battery is coupled to a dissipative environment, it can be observed that, under different battery dissipation rates, the stored energy of the quantum battery reaches its optimal value only when $\delta=1$ and $\Gamma_{C}=2g$. This demonstrates that the non-zero dissipation rate of the quantum battery does not affect its optimal charging conditions. In addition, when the temperature of the thermal reservoir and the dissipation rate of the battery are large, as shown in Fig. 3 (b), the stored energy of the battery does not change significantly with its dissipation rate.
@@
Figure 3: The steady-state stored energy of the quantum battery is shown as a function of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$. In panels (a) and (b), the temperature of the thermal reservoir is set to $T=0$ and $T=10$, respectively. The remaining parameters are the same as in Fig. 2.
@@
Figure 4: Under optimal charging conditions ($\Gamma_{C}=2g,\delta=1$), the stored energy of the quantum battery in steady state varies with the battery dissipation rate. Different curves represent different temperatures of the bosonic thermal reservoir. The green solid line, red dotted line, yellow dash-dotted line, and blue dashed line correspond to the bosonic thermal reservoir temperature $T$ of 0, 1, 10, and 50, respectively. The average thermal photon number as a function of temperature is given by $n_{b}=1/(e^{1/T}-1)$, where $T=0$ corresponds to $n_{b}=0$. All other parameters are the same as in Fig. 2.
@@
Under the perfect measurement ($\eta=1$) and the optimal charging conditions ($\Gamma_{C}=2g$, $\delta=1$), the stored energy of the battery becomes
@@
Based on Eq. (14), we explore the influence of quantum battery dissipation $\Gamma_{B}$ on optimal steady-state energy storage at different temperatures in Fig. 4, where each curve corresponds to different temperatures. It can be seen that when the dissipation rate $\Gamma_{B}$ of the battery is small, the optimal steady-state stored energy of the battery decreases with increasing temperature. When the dissipation rate $\Gamma_{B}$ of the battery gradually increases, the optimal stored energy of the battery increases and approaches a constant value as the temperature increases. It is easy to see from Eq. (14) that when $\Gamma_{B}\gg g$ and $n_{b}\rightarrow\infty$ the battery stored energy $E_{B}^{\rm optimal}(\infty)\approx 0.5\omega_{0}$.
Figure 5: The quantum battery is placed in a bosonic thermal reservoir: Panels (a) and (b) depict the ergotropy of the quantum battery as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger in the steady state. The environmental temperatures corresponding to (a) and (b) are $T=0$ and $T=10$, respectively. Solid red circles represent the optimal ergotropy of the battery, with the corresponding optimal charging parameters being $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. The other parameters are the same as in Fig. 2.
Figure 6: Panels (a) and (b) show the steady-state ergotropy of the quantum battery as a function of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$, with the thermal reservoir temperatures set to $T=0$ and $T=10$, respectively. The remaining parameters are identical to those in Fig. 2.
In addition to focusing on the stored energy of the quantum battery, the ergotropy of the quantum battery is also an important performance evaluation indicator. Similarly to Fig. 2, Fig. 5 plots the ergotropy of the quantum battery in the steady state as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger, where the dissipation rate of the battery is $\Gamma_{B}=0.1g$. The results demonstrate that optimal charging parameters still exist for the battery ergotropy. These optimal conditions coincide with those for battery energy storage, namely $\Gamma_{C}=2g$ and $\delta=1$. As shown in Fig. 6, the optimal charging conditions for the ergotropy of the quantum battery also remain independent of the dissipation rate of the battery $\Gamma_{B}$. Furthermore, when the dissipation rate $\Gamma_{B}$ of the quantum battery increases to a certain value, the ergotropy of the battery is zero regardless of the value of $\Gamma_{C}$ and $\delta$. This indicates that no energy can be extracted from the battery, and the feedback control mechanism becomes ineffective.
Under the optimal charging parameters ($\Gamma_{C}=2g,\delta=f/\Gamma_{C}=1$), the maximum extractable energy (the ergotropy) of the quantum battery in the bosonic thermal reservoir is
@@
Figure 7: The ergotropy of the quantum battery in the steady state is shown as a function of the battery dissipation rate $\Gamma_{B}$, where the dissipation rate of the charger is fixed at its optimal value, i.e., $\Gamma_{C}=2g$. The corresponding optimal feedback parameter $\delta$ is shown in the inset. Different curves correspond to different thermal reservoir temperatures. The green solid line, red dotted line, yellow dash-dotted line, and blue dashed line correspond to the bosonic thermal reservoir temperature $T$ of 0, 1, 10, and 50, respectively. The remaining parameters are the same as in Fig. 2.
@@
Figure 8: Under the optimal charging conditions in a bosonic thermal reservoir, the charging efficiency $R$ of the quantum battery is shown as a function of temperature $T$, with $\Gamma_{B}=0.1g$.
To evaluate the influence of the environment of the bosonic thermal reservoir on the battery charging efficiency $R$, we explored the impact of the temperature of the bosonic thermal reservoir on the quantum battery charging efficiency under optimal charging conditions ($\Gamma_{C}=2g,\delta=1$), as shown in Fig. 8, with $\Gamma_{B}=0.1g$. The definition of battery charging efficiency is given in Eq. (11). The results show that the charging efficiency $R$ of the quantum battery decreases with increasing temperature $T$, which means that the increase in the temperature of the reservoir inhibits the efficiency of energy extraction from the battery.
III.2 The charging process of quantum battery in fermionic thermal reservoir
Next, we consider the scenario in which the quantum battery is embedded in a finite-temperature fermionic thermal reservoir. By solving Eq. (7), we obtain that the stored energy of the quantum battery is given by
@@ III.2 The charging process of quantum battery in fermionic thermal reservoir
where we set $\delta=f/\Gamma_{C}$, $S=2\delta(\delta-\eta)+\eta$, $Q_{f}=\Gamma_{C}\delta^{2}+n_{f}\Gamma_{B}\eta$ and $W_{f}=\Gamma_{C}\Gamma_{B}(\Gamma_{C}+\Gamma_{B})[\Gamma_{B}\eta+\Gamma_{C}(2S-\eta)]$.
Figure 9: The quantum battery is placed in a fermionic reservoir: Panels (a) and (b) describe steady-state stored energy and the ergotropy of the quantum battery as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger, respectively. The temperature of the fermionic reservoir is $T=10$, the other parameters are the same as in Fig. 2.
For a zero-temperature thermal reservoir, the average excitation number of the thermal reservoir satisfies $n_{b(f)}=0$. As Eq. (7) shows, the master equation governing the dynamics of the control system has the same form for both bosonic and fermionic thermal reservoirs. Therefore, we will not specifically consider the scenario in which the temperature of the fermionic thermal reservoir is $T=0$. Figures 9 (a) and 9 (b) show the steady-state dependence of the battery’s stored energy and ergotropy on the feedback parameter $\delta$ and the charger dissipation rate $\Gamma_{C}$, respectively. Here, the temperature of the thermal reservoir is set at $T=10$ and the dissipation rate of the battery is set at $\Gamma_{B}=0.1g$. The results reveal the existence of optimal charging parameters that simultaneously maximize both the stored energy and the ergotropy of the quantum battery, which coincide with those obtained for the bosonic reservoir case. Furthermore, Fig. 10 again confirms that these parameters remain independent of the battery dissipation rate $\Gamma_{B}$.
Figure 10: Panels (a) and (b) show the steady-state stored energy and ergotropy of the quantum battery, respectively, as functions of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$. The fermionic reservoir temperature is set to $T=10$. All other parameters are the same as in Fig. 2.
Figure 11: Under optimal charging conditions ($\Gamma_{C}=2g,\delta=1$), the steady-state stored energy of the quantum battery is shown as a function of the battery dissipation rate. Different curves correspond to different reservoir temperatures. The green solid line, blue dashed line, red dotted line, and yellow dash-dotted line correspond to the fermionic thermal reservoir temperature $T$ of 0, 0.5, 1, and 10, respectively. The average thermal excitation number follows $n_{f}=1/(e^{1/T}+1)$, where $T=0$ corresponds to $n_{f}=0$. Other parameters match those in Fig. 2.
Under the perfect measurement ($\eta=1$) and the optimal charging conditions ($\Gamma_{C}=2g$, $\delta=1$), the stored energy of the battery becomes
@@
Figure 12: The steady-state ergotropy of the quantum battery is shown as a function of $\Gamma_{B}$, with different curves corresponding to different fermionic thermal reservoir temperatures. The dissipation rate of the charger is set to $\Gamma_{C}=2g$, and the corresponding optimal feedback parameters $\delta$ are shown in the inset.
Under optimal charging parameters ($\Gamma_{C}=2g$, $\delta=1$), the maximum extractable energy (the ergotropy) of the quantum battery can be expressed as:
$\mathcal{E}(\infty)$ $=$ $\omega_{0}\frac{\beta_{f}+\lvert\beta_{f}\rvert}{(2g+\Gamma_{B})^{2}},$ (19)
@@
$\mathcal{E}(\infty)$ $=$ $\omega_{0}\frac{2\beta_{f}}{(2g+\Gamma_{B})^{2}}$ (20)
$=$ $\omega_{0}\left[\frac{8g\Gamma_{B}+2\Gamma_{B}^{2}}{(2g+\Gamma_{B})^{2}}n_{f}+(4g^{2}-4g\Gamma_{B}-\Gamma_{B}^{2})\right].$
Figure 13: Under the optimal charging conditions of the fermionic thermal reservoir, the charging efficiency $R$ of the quantum battery is a function of the temperature $T$, with $\Gamma_{B}=0.1g$.
The above equation indicates that, for fixed $g$ and $\Gamma_{B}$, the extractable energy from the quantum battery increases with temperature. In summary, when the quantum battery is coupled to a fermionic thermal reservoir, higher temperatures not only broaden the permissible range of the battery’s dissipation rate but also enhance the battery’s energy extraction capability.
Figure 14: The multi-particle quantum battery is placed in a bosonic thermal reservoir, its optimal energy storage is analyzed as a function of the number of constituent quantum battery particles at different reservoir temperatures. The coupling strengths between the internal particles of the quantum battery in panels (a), (b), and (c) are $J=0$, $J=0.1g$, and $J=g$, respectively. The corresponding optimal charger dissipation parameters $\Gamma_{C}$ are shown in panels (d), (e), and (f). The remaining parameters are set as $g=0.01\omega_{0}$, $\delta=f/\Gamma_{C}=1$, and $\Gamma_{B}=0.05g$.
Figure 12 illustrates the dependence of the quantum battery ergotropy on its dissipation rate, where the dissipation rate of the charger remains at its optimal value ($\Gamma_{C}=2g$). The corresponding optimal feedback parameters are presented in the inset. The results demonstrate that increasing temperature promotes the extraction of battery energy. Similarly to the bosonic thermal reservoir case, there exists a critical value for the dissipation rate $\Gamma_{B}$ of the battery in terms of energy extraction. In particular, when the battery is coupled to a fermionic thermal reservoir, an increase in temperature extends the critical point of the dissipation rate of the battery $\Gamma_{B}$ and improves its usability. Subsequently, we explore the influence of the temperature of the fermionic thermal reservoir on the charging efficiency $R$ of the battery under optimal charging conditions through Fig. 13, where the dissipation rate of the battery is $\Gamma_{B}=0.1g$. The results show that the charging efficiency $R$ increases with temperature, suggesting that the temperature can serve as a resource to improve battery performance. This is different from the scheme proposed in reference Song et al. (2024b), which exhibits the temperature will be unfavorable for the extraction of battery energy, when the environment is within the positive temperature range, while in our scheme, temperature is still a favorable factor for battery charging. We have conducted a simple comparative analysis of these two schemes in the Appendix B.

### v3

@@ II.2 Related performance parameter of the quantum battery
where $H_{B}=\omega_{0}\sigma_{B}^{+}\sigma_{B}^{-}$, $\rho_{B}(t)={\rm Tr}_{C}[\rho(t)]$, and $\rho_{B}(0)=|g\rangle_{B}\langle g|$ are the free Hamiltonian, the reduced density matrix, and the initial state of the quantum battery, respectively. When the quantum battery is in the excited state $|e\rangle$, it means that it is fully charged, and the corresponding maximum stored energy is $E_{B}^{\rm max}=\omega_{0}$. To evaluate the useful work of the quantum battery storage, we use the concept of ergotropy Allahverdyan et al. (2004); Francica et al. (2017) and its form is
@@
which describes the maximum energy that can be extracted from the battery. The second term in Eq. (9) indicates the minimum energy that cannot be extracted from the battery by executing all unitaries $U$ in the system, and the state corresponding to the minimum energy is known as the passive state Pusz and Woronowicz (1978); Lenard (1978).
@@
III CHARGING PROCESS OF A QUANTUM BATTERY IN DIFFERENT THERMAL RESERVOIR ENVIRONMENTS
III.1 Charging process of quantum battery in a bosonic thermal reservoir
In this section, we first consider the case where the quantum battery is in a bosonic thermal reservoir. After setting $\mathop{\dot{\rho}}=0$ in Eq. (7), the steady-state solutions of the system are obtained (see Appendix A). Combined with Eq. (8), the stored energy of the battery is finally obtained as
@@ III.1 Charging process of quantum battery in a bosonic thermal reservoir
This clearly shows that when $\delta=f/\Gamma_{C}=1$, the stored energy of the battery reaches its maximum value, that is, $E_{B}^{\rm max}(\infty)=\omega_{0}$. This implies that the effect of spontaneous emission on the charger can be offset by adjusting the feedback intensity $f$, enabling full charging of the quantum battery, this result is consistent with Ref. Mitchison et al. (2021). However, in practice, a quantum battery cannot be perfectly isolated from its surroundings, so it is necessary to consider placing it in an open system. To discuss the influence of feedback strength $f$ and charger dissipation on quantum battery energy storage, we conduct corresponding numerical simulations according to Eq. (12). In the absence of special instructions, the remaining parameters are set as $\eta=1$, $g=0.01\omega_{0}$, and $\Gamma_{B}=0.1g$.
@@
Figure 2: Under a bosonic thermal reservoir, the stored energy of the quantum battery is a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger. The temperature of the bosonic thermal reservoir is (a) $T=0$ $(n_{b}=0)$ and (b)$T=10$ ($n_{b}\approx 9.51$). Red circles represent the optimal stored energy of the quantum battery, with the corresponding optimal charging parameters being $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. The other parameters are $g=0.01\omega_{0}$ and $\Gamma_{B}=0.1g$, respectively.
@@
Figures 2(a) and 2(b) show the stored energy of the quantum battery as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger when the battery is in a dissipative bosonic thermal reservoir with temperatures $T=0$ ($n_{b}=0$) and $T=10$ ($n_{b}\approx 9.51$), respectively. The results show that the optimal stored energy of the battery is regulated by the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger. There exists a set of optimal parameters $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. Here $\delta=1$ means that the feedback strength $f$ is equal to the dissipation rate $\Gamma_{C}$ of the charger (i.e.,$f=\Gamma_{C}$).
@@
To determine whether optimal charging conditions are affected by the dissipation rate of the battery $\Gamma_{B}$, we explore the influence of the feedback parameter $\delta$ and the dissipation rate of the charger and the quantum battery on the stored energy of the quantum battery in steady state in Fig. 3. The thermal reservoir temperatures for Figs. 3(a) and 3(b) are $T=0$ ($n_{b}=0$) and $T=10$ ($n_{b}\approx 9.51$), respectively. The results show that when the battery dissipation rate is zero ($\Gamma_{B}=0$), there is no restriction on the magnitude of the dissipation rate $\Gamma_{C}$ of the charger; as long as the optimal feedback parameter satisfies $\delta=f/\Gamma_{C}=1$, the quantum battery can be fully charged. This is consistent with the results of the previous analysis. However, when the quantum battery is coupled to a dissipative environment, it can be observed that, under different battery dissipation rates, the stored energy of the quantum battery reaches its optimal value only when $\delta=1$ and $\Gamma_{C}=2g$. This demonstrates that the nonzero dissipation rate of the quantum battery does not affect its optimal charging conditions. In addition, when the temperature of the thermal reservoir and the dissipation rate of the battery are high, as shown in Fig. 3(b), the stored energy of the battery does not change significantly with its dissipation rate.
@@
Figure 3: Steady-state stored energy of the quantum battery plotted as a function of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$. The temperature of the thermal reservoir is set to (a) $T=0$ and (b) $T=10$. The remaining parameters are the same as in Fig. 2.
@@
Figure 4: Under optimal charging conditions $\Gamma_{C}=2g$ and $\delta=1$, the stored energy of the quantum battery in steady state varies with the battery dissipation rate. Different curves represent different temperatures of the bosonic thermal reservoir. The green solid line, red dotted line, yellow dash-dotted line, and blue dashed line correspond to the bosonic thermal reservoir temperatures $T$ of 0, 1, 10, and 50, respectively. The average thermal photon number as a function of temperature is given by $n_{b}=1/(e^{1/T}-1)$, where $T=0$ corresponds to $n_{b}=0$. All other parameters are the same as in Fig. 2.
@@
Under the perfect measurement ($\eta=1$) and the optimal charging conditions ($\Gamma_{C}=2g$ and $\delta=1$), the stored energy of the battery becomes
@@
Based on Eq. (14), we explore the influence of quantum battery dissipation $\Gamma_{B}$ on optimal steady-state energy storage at different temperatures in Fig. 4, where each curve corresponds to a different temperature. It can be seen that when the dissipation rate $\Gamma_{B}$ of the battery is low, the optimal steady-state stored energy of the battery decreases with increasing temperature. When the dissipation rate $\Gamma_{B}$ of the battery gradually increases, the optimal stored energy of the battery increases and approaches a constant value as the temperature increases. It is easy to see from Eq. (14) that when $\Gamma_{B}\gg g$ and $n_{b}\rightarrow\infty$ the battery stored energy $E_{B}^{\rm optimal}(\infty)\approx 0.5\omega_{0}$.
Figure 5: The quantum battery is placed in a bosonic thermal reservoir. The ergotropy of the quantum battery is depicted as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger in the steady state for environmental temperatures corresponding to (a) $T=0$ and (b) $T=10$. Red circles represent the optimal ergotropy of the battery, with the corresponding optimal charging parameters being $\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$. The other parameters are the same as in Fig. 2.
Figure 6: Steady-state ergotropy of the quantum battery as a function of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$, with the thermal reservoir temperatures set to (a) $T=0$ and (b) $T=10$. The remaining parameters are identical to those in Fig. 2.
In addition to focusing on the stored energy of the quantum battery, the ergotropy of the quantum battery is also an important performance evaluation indicator. Similarly to Fig. 2, Fig. 5 plots the ergotropy of the quantum battery in the steady state as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger, where the dissipation rate of the battery is $\Gamma_{B}=0.1g$. The results demonstrate that optimal charging parameters still exist for the battery ergotropy. These optimal conditions coincide with those for battery energy storage, namely, $\Gamma_{C}=2g$ and $\delta=1$. As shown in Fig. 6, the optimal charging conditions for the ergotropy of the quantum battery also remain independent of the dissipation rate of the battery $\Gamma_{B}$. Furthermore, when the dissipation rate $\Gamma_{B}$ of the quantum battery increases to a certain value, the ergotropy of the battery is zero regardless of the value of $\Gamma_{C}$ and $\delta$. This indicates that no energy can be extracted from the battery, and the feedback control mechanism becomes ineffective.
Under the optimal charging parameters ($\Gamma_{C}=2g$ and $\delta=f/\Gamma_{C}=1$), the maximum extractable energy (the ergotropy) of the quantum battery in the bosonic thermal reservoir is
@@
Figure 7: Ergotropy of the quantum battery in the steady state plotted as a function of the battery dissipation rate $\Gamma_{B}$, where the dissipation rate of the charger is fixed at its optimal value, i.e., $\Gamma_{C}=2g$. The corresponding optimal feedback parameter $\delta$ is shown in the inset. Different curves correspond to different thermal reservoir temperatures. The green solid line, red dotted line, yellow dash-dotted line, and blue dashed line correspond to bosonic thermal reservoir temperatures $T$ of 0, 1, 10, and 50, respectively. The remaining parameters are the same as in Fig. 2.
@@
Figure 8: Under the optimal charging conditions in a bosonic thermal reservoir, the charging efficiency $R$ of the quantum battery is plotted as a function of temperature $T$, with $\Gamma_{B}=0.1g$.
To evaluate the influence of the environment of the bosonic thermal reservoir on the battery charging efficiency $R$, we explore the impact of the temperature of the bosonic thermal reservoir on the quantum battery charging efficiency under optimal charging conditions ($\Gamma_{C}=2g$ and $\delta=1$), as shown in Fig. 8, with $\Gamma_{B}=0.1g$. The definition of battery charging efficiency is given in Eq. (11). The results show that the charging efficiency $R$ of the quantum battery decreases with increasing temperature $T$, which means that the increase in the temperature of the reservoir inhibits the efficiency of energy extraction from the battery.
III.2 Charging process of quantum battery in a fermionic thermal reservoir
Next we consider the scenario in which the quantum battery is embedded in a finite-temperature fermionic thermal reservoir. By solving Eq. (7), we obtain that the stored energy of the quantum battery is given by
@@ III.2 Charging process of quantum battery in a fermionic thermal reservoir
where we set $\delta=f/\Gamma_{C}$, $S=2\delta(\delta-\eta)+\eta$, $Q_{f}=\Gamma_{C}\delta^{2}+n_{f}\Gamma_{B}\eta$, and $W_{f}=\Gamma_{C}\Gamma_{B}(\Gamma_{C}+\Gamma_{B})[\Gamma_{B}\eta+\Gamma_{C}(2S-\eta)]$.
Figure 9: The quantum battery is placed in a fermionic reservoir. Plotted are (a) the steady-state stored energy and (b) the ergotropy of the quantum battery, as a function of the feedback parameter $\delta$ and the dissipation rate $\Gamma_{C}$ of the charger. The temperature of the fermionic reservoir is $T=10$ and the other parameters are the same as in Fig. 2.
For a zero-temperature thermal reservoir, the average excitation number of the thermal reservoir satisfies $n_{b(f)}=0$. As Eq. (7) shows, the master equation governing the dynamics of the control system has the same form for both bosonic and fermionic thermal reservoirs. Therefore, we will not specifically consider the scenario in which the temperature of the fermionic thermal reservoir is $T=0$. Figures 9(a) and 9(b) show the steady-state dependence of the battery’s stored energy and ergotropy on the feedback parameter $\delta$ and the charger dissipation rate $\Gamma_{C}$, respectively. Here the temperature of the thermal reservoir is set at $T=10$ and the dissipation rate of the battery is set at $\Gamma_{B}=0.1g$. The results reveal the existence of optimal charging parameters that simultaneously maximize both the stored energy and the ergotropy of the quantum battery, which coincide with those obtained for the bosonic reservoir case. Furthermore, Fig. 10 again confirms that these parameters remain independent of the battery dissipation rate $\Gamma_{B}$.
Figure 10: (a) Steady-state stored energy and (b) ergotropy of the quantum battery, as a function of $\delta$, $\Gamma_{B}$, and $\Gamma_{C}$. The fermionic reservoir temperature is set to $T=10$. All other parameters are the same as in Fig. 2.
Figure 11: Under optimal charging conditions $\Gamma_{C}=2g$ and $\delta=1$, the steady-state stored energy of the quantum battery is plotted as a function of the battery dissipation rate. Different curves correspond to different reservoir temperatures. The green solid line, blue dashed line, red dotted line, and yellow dash-dotted line correspond to the fermionic thermal reservoir temperatures $T$ of 0, 0.5, 1, and 10, respectively. The average thermal excitation number follows $n_{f}=1/(e^{1/T}+1)$, where $T=0$ corresponds to $n_{f}=0$. The other parameters match those in Fig. 2.
Under the perfect measurement ($\eta=1$) and the optimal charging conditions ($\Gamma_{C}=2g$ and $\delta=1$), the stored energy of the battery becomes
@@
Figure 12: Steady-state ergotropy of the quantum battery plotted as a function of $\Gamma_{B}$, with different curves corresponding to different fermionic thermal reservoir temperatures. The dissipation rate of the charger is set to $\Gamma_{C}=2g$ and the corresponding optimal feedback parameters $\delta$ are shown in the inset.
Under optimal charging parameters ($\Gamma_{C}=2g$ and $\delta=1$), the maximum extractable energy (the ergotropy) of the quantum battery can be expressed as
$\mathcal{E}(\infty)$ $=$ $\omega_{0}\frac{\beta_{f}+\lvert\beta_{f}\rvert}{2(2g+\Gamma_{B})^{2}},$ (19)
@@
$\mathcal{E}(\infty)$ $=$ $\omega_{0}\frac{2\beta_{f}}{2(2g+\Gamma_{B})^{2}}$ (20)
$=$ $\omega_{0}\left[\frac{8g\Gamma_{B}+2\Gamma_{B}^{2}}{(2g+\Gamma_{B})^{2}}n_{f}+\frac{4g^{2}-4g\Gamma_{B}-\Gamma_{B}^{2}}{(2g+\Gamma_{B})^{2}}\right].$
Figure 13: Under the optimal charging conditions of the fermionic thermal reservoir, the charging efficiency $R$ of the quantum battery is plotted as a function of the temperature $T$, with $\Gamma_{B}=0.1g$.
This equation indicates that, for fixed $g$ and $\Gamma_{B}$, the extractable energy from the quantum battery increases with temperature. In summary, when the quantum battery is coupled to a fermionic thermal reservoir, higher temperatures not only broaden the permissible range of the battery’s dissipation rate but also enhance the battery’s energy extraction capability.
Figure 14: The multiparticle quantum battery is placed in a bosonic thermal reservoir, and its optimal energy storage is analyzed as a function of the number of constituent quantum battery particles at different reservoir temperatures. The coupling strengths between the internal particles of the quantum battery are (a) $J=0$, (b) $J=0.1g$, and (c) $J=g$. The corresponding optimal charger dissipation parameters $\Gamma_{C}$ are shown in (d)-(f). The remaining parameters are set as $g=0.01\omega_{0}$, $\delta=f/\Gamma_{C}=1$, and $\Gamma_{B}=0.05g$.
Figure 12 illustrates the dependence of the quantum battery ergotropy on its dissipation rate, where the dissipation rate of the charger remains at its optimal value ($\Gamma_{C}=2g$). The corresponding optimal feedback parameters are presented in the inset. The results demonstrate that increasing temperature promotes the extraction of battery energy. Similarly to the bosonic thermal reservoir case, there exists a critical value for the dissipation rate $\Gamma_{B}$ of the battery in terms of energy extraction. In particular, when the battery is coupled to a fermionic thermal reservoir, an increase in temperature extends the critical point of the dissipation rate of the battery $\Gamma_{B}$ and improves its usability. Subsequently, we explore the influence of the temperature of the fermionic thermal reservoir on the charging efficiency $R$ of the battery under optimal charging conditions through Fig. 13, where the dissipation rate of the battery is $\Gamma_{B}=0.1g$. The results show that the charging efficiency $R$ increases with temperature, suggesting that the temperature can serve as a resource to improve battery performance. This is different from the scheme proposed in Ref. Song et al. (2024b), which exhibits the temperature will be unfavorable for the extraction of battery energy, when the environment is within the positive temperature range, while in our scheme, temperature is still a favorable factor for battery charging. We conduct a simple comparative analysis of these two schemes in Appendix B.